OpenStudy (anonymous):

Determine the average rate of change of the function between the given values of the variable. f(t)=2/t; t=a, t=a+h I get (2a-2)/(-a)

OpenStudy (amistre64):

the average rate of change is usually defined as :\[\frac{f(x+h)=f(x)}{(x+h)-(x)}\]

OpenStudy (amistre64):

*= meant to be a -

OpenStudy (anonymous):

I'm using that, not sure where I'm going wrong through

OpenStudy (amistre64):

\[\frac{\frac{2}{a+h}-\frac{2}{a}}{a+h-a}\] \[\frac{\frac{2a-2(a+h)}{a}}{h}\] \[\frac{\frac{2a-2a-2h}{a}}{h}\] \[\frac{\frac{-2h}{a}}{h}\] \[\frac{-2\cancel{h}}{a\cancel{h}}=-\frac{2}{a}\] if i did it right

OpenStudy (amistre64):

i see a place i messed it at

OpenStudy (anonymous):

the answer key says (-2)/a(a+h)

OpenStudy (amistre64):

i shoulda combined my denoms :) \[\frac{\frac{2a-2(a+h)}{a(a+h)}}{h}\]

OpenStudy (amistre64):

but yeah, it works out to that

OpenStudy (amistre64):

i wouldnt know where your going wrong at either,since i cant see your work :)

OpenStudy (anonymous):

i got to the first part of your post but am kind of confused at the second

OpenStudy (amistre64):

the second part is just working thru the math .... which part gets you lost? other than my mess up with adding the fractions :/